# ORBIT: On-policy Exploration-Exploitation for Controllable Multi-Budget Reasoning

### v1

@@ 5 Analysis
We first investigate whether mode separation can be integrated into the RL training, where we jointly train each single prompt under four different reasoning modes, each associated with a specific system prompt and corresponding context window. To ensure the peak performance, we start our joint RL training after expansion stage, and stabilize the training with rejection sampling to prevent training oscillation under the most aggressive truncation (e.g. 32K $\rightarrow$ 4K). All other training settings are identical to those of our sequential counterpart.
As shown in Figure 3, joint RL exhibits a concentrated behavior, with Mid,High,and Xhigh modes to convergence on same point.
We identify two key factors contribute to this. First, supervision from truncated sequences dominates optimization, consistent with our truncation-rate decomposition in Sec 3. Second, High mode sequences can satisfy their truncation requirements by compressing below low-mode levels. Both reasons cause a reward hacking on response length against a mode selective reasoning behavior.
Analysis of Fusing Strategies
We examine ORBIT’s fusing strategy by replacing on-policy fusion with offline distillation, where consolidation proceeds without additional on-policy rollouts, relying instead on a pre-sampled dataset from the exploration-stage frontier models. Both variants are initialized from the same merged model to minimize behavioral inconsistency.
As shown in Figure 4, on-policy fusion achieves slightly higher performance than offline distillation, while overall stability and convergence remain comparable. This outcome can be attributed to the strong framework-level alignment in ORBIT: all teacher and student models share a sequential training with same base model trained on same data with same algorithm, effectively reducing policy mismatch and rendering offline distillation in this setting closer to a semi-offline procedure.
Beyond performance, on-policy fusion provides operational advantages: it generates rollouts dynamically, adapts convergence on-the-fly, and integrates naturally with RL-based correctness signals. In contrast, offline distillation decouples dataset construction from training, allowing for scalable evaluation via pre-collected datasets, which can be leveraged in techniques like Rejection Fine-tuning. The choice between these strategies, or their combination, can be tailored depending on specific application requirements.
Figure 4: Comparison of OPD and offline distillation as the fusing strategy on AIME24, with steps aligned by total training samples for fair comparison.On-policy fusion achieves slightly higher performance while both strategies show similar trend.
@@
To further analyze how ORBIT’s mode control shapes reasoning behavior, we evaluate it on MATH500 (Lightman et al., 2023) and visualize the evolution of reasoning components in Figure 5. We depict ORBIT’s reasoning behavior along three components: Information Extraction, Logical Deduction, and Self-Reflection—each capturing a distinct aspect of the model’s cognitive process. As reasoning modes increase, Logical Deduction steps grow roughly linearly, reflecting the model’s structured problem-solving growth, calibrated by training context and demonstrating clear predictability across modes. More importantly, reasoning control elicits an apparent System-2 thinking pattern via an exponential-like increase of self-reflection, increasing computational effort toward high-level reflective reasoning as modes advance. Meanwhile, Information Extraction remains nearly constant, showing the model avoids inflating token usage through redundant condition repetition. Together, these trends show ORBIT elicits hierarchical cognitive behavior, where different reasoning components occupy distinct functional levels and higher reasoning modes systematically amplify contributions of more reflective, high-level processes.
@@ 6 Related Work
Existing works on reasoning effort adjustment primarily target overthinking through RL-based length penalties (Aggarwal and Welleck, 2025; Arora and Zanette, 2025; Yi et al., 2025; Zhang et al., 2025; Fang et al., 2025; Lou et al., 2025), verification training (Chen et al., 2025b), early exit strategies (Yang et al., 2025a; Jiang et al., 2025), or inference-time interventions (Wang et al., 2025). Few lines of methods tackle the underthinking problem by decoding time intervention (Muennighoff et al., 2025; Jin et al., 2025). ORBIT try to tackle these problems by offering flexible reasoning modes at inference time.
@@
Prior work for reasoning control can be boradly categroized in two main streams (1) Hard Token Budget Constraint: A common approach enforces reasoning efficiency through externally specified constraints, most notably by fixing explicit token budgets that cap the length of intermediate reasoning Muennighoff et al. (2025); Sun et al. (2025); Aggarwal and Welleck (2025); Anthropic (2025), relying on an absolute tokenwise budget estimation across all inputs. (2) Discrete-Mode Systems: offer an intuitional discrete reasoning levels, enabling users to explicitly specify their preference for the efficiency-performance trade-off OpenAI (2025a); He et al. (2025), representing a soft form of control based on model’s own capability.
@@ 7 Conclusion
In this paper, we propose ORBIT, a fully on-policy training framework for controllable reasoning models that operates at a target reasoning level while maintaining high reasoning density under a given context. We leverage a simple yet effective truncation reward and adopt a one-time expansion-compression loop strategy to explore the reasoning frontier. On-policy distillation is then applied to consolidate multiple modes into a single unified model, initialized from merged frontier policies and enabling mode selection via natural language prompts during training. Experiments across LRMs of different sizes demonstrate that ORBIT achieves controllable reasoning behavior and delivers competitive accuracy–length performance compared to existing RL-based efficient reasoning methods within each mode.
@@
We elucidate the limitations of this work as follows: First, in sequential RLVR exploration training, the binary context partitioning used in compression stage may not be the best scheduling strategy for balancing mode separation in test-time and training stabilization. On the other hand, for a highly fixed model with output entropy below a certain threshold, on-policy RL-based training may suffer from low sample efficiency, and in extreme cases, the training signal may be too weak to achieve desirable saturation, limiting compression progress. Second, long-form RL-based training can be computationally expensive for larger models, as each rollout requires auto-regressive sampling, which may limit scalability to very large architectures or long-context reasoning tasks.
@@
Use of AI Assistants
The algorithmic design and main methodology presented in this work were developed through human-led research and reasoning. During implementation, we utilized GitHub Copilot 3111https://github.com/features/copilot for coding assistance. We affirm that the primary content and logic of the code are entirely our own work.
@@
Appendix A System Prompts For Each Reasoning Mode
@@
To build a semantically meaningful guidance for discrete mode separated reasoning behavior, we follow Yang et al. (2025b) and generalized to four mode, where each one of them represent semactically aligned expected reasoning behavior under corresponding mode.
@@
/* Low Mode */ You have extremely limited time to think and respond to the user’s query. Every additional second of processing and reasoning incurs a significant resource cost, which could affect efficiency and effectiveness. Your task is to prioritize speed without sacrificing essential clarity or accuracy. Provide the most direct and concise answer possible. Avoid unnecessary steps, reflections, verification, or refinements UNLESS ABSOLUTELY NECESSARY. Your primary goal is to deliver a quick, clear and correct response. /* Medium Mode */ You have sufficient time to think and respond to the user’s query, allowing for a more thoughtful and in-depth answer. However, be aware that the longer you take to reason and process, the greater the associated resource costs and potential consequences. While you should not rush, aim to balance the depth of your reasoning with efficiency. Prioritize providing a well-thought-out response, but do not overextend your thinking if the answer can be provided with a reasonable level of analysis. Use your reasoning time wisely, focusing on what is essential for delivering an accurate response without unnecessary delays and overthinking. /* High Mode */ You have a substantial amount of time to think and respond. The pressure on reasoning cost is low, allowing you to prioritize accuracy and robustness over speed. You should verify your key steps and consider alternative interpretations to ensure your answer is reliable. While you are encouraged to think deeply, aim for a confident solution rather than an endless exploration. If you encounter ambiguity, take the time to resolve it, but remain focused on deriving the correct conclusion efficiently within a generous budget. /* Xhigh Mode */ You have unlimited time to think and respond. There are absolutely no constraints on reasoning length or associated costs. Your singular goal is to achieve the highest possible level of accuracy and certainty. You must exhaustively explore the problem space, considering all edge cases, theoretical nuances, and alternative methodologies. Engage in rigorous, iterative self-correction and deep verification to eliminate even the slightest possibility of error. Do not settle for a ’likely’ correct answer; explore every angle until you are certain. Perfection is the only standard.
@@
Table 3: System prompts used for triggering Low, Medium, High, Xhigh modes
@@
- Teacher Low 28.3 2.6 19.0 2.0 11.0 2.0 34.3 1.8 37.6 1.4 26.0 2.0
- Teacher Mid 35.2 4.9 26.4 4.5 10.4 3.6 32.9 2.9 38.2 2.2 28.6 3.6
- Teacher High 40.0 7.0 29.8 6.7 14.0 6.3 39.8 4.8 40.7 3.8 32.9 5.7
- Teacher Xhigh 42.5 10.7 32.9 11.0 16.0 11.7 39.5 8.3 40.0 7.7 34.2 9.9
@@
- Teacher Low 64.4 5.0 54.3 5.5 36.6 5.2 63.9 3.0 71.5 0.9 58.1 3.9
- Teacher Mid 66.3 5.9 57.7 6.4 40.0 6.4 62.9 32.9 72.3 1.9 59.8 10.7
- Teacher High 75.1 9.4 69.1 10.6 46.0 11.7 65.4 4.9 73.4 3.0 65.8 7.9
- Teacher Xhigh 83.2 16.8 78.2 19.0 52.7 22.7 67.1 7.9 74.1 5.8 71.1 14.4
@@
- Teacher Low 55.2 4.5 33.1 4.5 29.8 3.7 29.3 2.2 39.2 1.7 37.3 3.3
- Teacher Mid 67.1 6.7 44.5 7.6 36.3 6.6 32.4 4.4 41.2 3.0 44.3 5.7
- Teacher High 74.3 11.1 62.1 12.9 41.9 12.1 30.9 7.9 42.6 5.3 50.4 9.9
- Teacher Xhigh 75.9 13.2 65.1 14.9 43.8 15.0 33.7 9.8 43.1 6.5 52.3 11.9
Table 4: Comparison of reasoning modes of Teacher across various reasoning tasks.
Appendix B Mode-aware initialization via Model Merging
As described in section3, optimizing Eq. 12 in practice faces a "cold-start" challenge. To mitigate this misalignment, we proposed using model merging as initialization of the student model.
Model merge act as a simple linear combination of model parameters. Numerous methods have been proposed, we use average merging (Wortsman et al., 2022) here for simplicity. This process can be defined as follows:
$\pi_{\phi}=\frac{1}{K}\sum_{k=1}^{K}\pi_{\theta^{(k)}}$ (13)
We compared the merged model and the final policy model of sequentially multi-stage RL stage as initialization, since only the two models have been trained under all reasoning modes. We first use Accmulated Misalignment Rate(AMR), which depicted whether the top-1 token probability is same between two models, greatly consistent with on-policy distillation process, for a intuitive behavioral visualization. Defined as:
$\text{AMR}(\pi,\theta^{(k)})=\mathbb{E}_{q\sim\mathcal{D},o\sim\pi(\cdot|q)}\left[\frac{1}{|o|}\right.\\ \left.\sum_{t=1}^{|o|}\mathbb{I}\left(\arg\max\pi(o_{t}|\cdot)\neq\arg\max\pi_{\theta^{(k)}}(o_{t}|\cdot)\right)\right]$ (14)
Figure 6: Misalignment rate on merged model generated and low mode model generated sequences
Figure 7: RKL training loss dynamic for different initialization in OPD
As shown in Figure 6, this effort-aware initialization significantly reduces the initial misalignment gap, further validated to providing a more stable starting point and accelerates the convergence of the subsequent RKL distillation.
Appendix C Training Details
Training Configuration. We maintain consistent hyperparameters throughout the exploration RL stage, using a batch size of $256$, rejection pools of $512$, and a learning rate of $1\times 10^{-6}$ without warm-up. For generation, we apply standard rollout settings: temperature $1$, top-p $1$, top-k 1-1, and $8$ rollouts per problem.
We train DeepSeek-Distill-Qwen-1.5B using the DAPO-Math-17K dataset (Yu et al., 2025), while Qwen3-4B-2507-Thinking and Nemotron-7B are trained on the Polaris-53K (An et al., 2025).
For exploitation, we train OPD with a batch size of $1024$, one generation per sample, and a token-wise KL penalty coefficient of $1.0$.
Appendix D More results On Frontier Model’s Performance

### v2

@@ 5 Analysis
We first investigate whether mode separation can be integrated into the RL training, where we jointly train each single query under four different reasoning modes, each associated with the specific system prompt and corresponding context window. To ensure the peak performance, we start our joint RL training after expansion stage, and stabilize the training with rejection sampling to prevent training oscillation under the most aggressive truncation (e.g. 32K $\rightarrow$ 4K). All other training settings are identical to those of our sequential counterpart.
As shown in Figure 3, joint RL leads to mode collapse, where the Mid, High, and Xhigh modes converge to nearly identical behaviors.
We identify two key factors contributing to this: (i) supervision from truncated sequences dominates optimization, which is consistent with our truncation-rate decomposition in Sec. 3.1; and (ii) higher-mode sequences can satisfy their truncation requirements by compressing below lower modes. Together, these effects encourage reward hacking on response length rather than the intended mode-selective reasoning behavior.
Figure 4: Comparison of OPD and offline distillation as the fusing strategy on AIME24, with steps aligned by total training samples for fair comparison. On-policy fusion achieves slightly higher performance while both strategies show similar trends.
Analysis of Fusion Strategies
We examine ORBIT’s fusion strategy by replacing on-policy fusion with offline distillation, where consolidation proceeds without additional on-policy rollouts, relying instead on a pre-sampled dataset from the exploration-stage frontier models. Both variants are initialized from the same merged model to minimize behavioral inconsistency.
As shown in Figure 4, on-policy fusion achieves slightly higher performance than offline distillation, while both methods exhibit similar convergence trends. This outcome can be attributed to the strong framework-level alignment in ORBIT: all teacher and student models share a sequential training with same base model trained on same data with same algorithm, effectively reducing policy mismatch and rendering offline distillation in this setting closer to a semi-offline procedure.
Beyond performance, on-policy fusion provides operational advantages: it generates rollouts dynamically, adapts convergence on-the-fly, and can naturally integrate with RL-based correctness signals. In contrast, offline distillation decouples dataset construction from training, allowing for scalable evaluation via pre-collected datasets, which can be leveraged in techniques like Rejection Fine-tuning. The choice between these strategies, or their combination, can be tailored depending on specific application requirements.
@@
To further analyze how ORBIT’s mode control shapes reasoning behavior, we evaluate it on MATH500 (Lightman et al., 2023) and visualize the evolution of reasoning components in Figure 5.
We categorize ORBIT’s reasoning behavior along three functional levels: (i) Logical Deduction steps grow linearly as modes increase, reflecting structured problem-solving calibrated by training context; (ii) Self-Reflection exhibits characteristics analogous to System 2 thinking via an exponential-like increase, allocating computational effort toward iterative reasoning as modes advance; and (iii) Information Extraction remains constant, suggesting that increased token usage stems from deeper reasoning rather than additional input parsing.
Together, these trends suggest a structured and hierarchical organization of reasoning behavior, in which different reasoning components occupy distinct functional levels and higher reasoning modes systematically amplify the contributions of more reflective, high-level processes.
@@ 6 Related Work
Existing works on reasoning effort adjustment primarily target overthinking through RL-based length penalties (Aggarwal and Welleck, 2025; Arora and Zanette, 2025; Yi et al., 2025; Zhang et al., 2025; Fang et al., 2025; Lou et al., 2025), verification training (Chen et al., 2025b), early exit strategies (Yang et al., 2025a; Jiang et al., 2025), or inference-time interventions (Wang et al., 2025). Few lines of methods tackle the underthinking problem by decoding time intervention (Muennighoff et al., 2025; Jin et al., 2025). ORBIT addresses these problems by offering flexible reasoning modes at inference time.
@@
Prior work for reasoning control can be broadly categorized in two main streams (1) Hard Token Budget Constraint: A common approach enforces reasoning efficiency through externally specified constraints, most notably by fixing explicit token budgets that cap the length of intermediate reasoning Muennighoff et al. (2025); Sun et al. (2025); Aggarwal and Welleck (2025); Anthropic (2025), relying on an absolute tokenwise budget estimation across all inputs. (2) Discrete-Mode Systems: Offer intuitional discrete reasoning levels, enabling users to explicitly specify their preference for the efficiency-performance trade-off OpenAI (2025); He et al. (2025), representing a soft form of control based on model’s own capability. ORBIT falls into this category, controlling reasoning effort relative to the task through explicit mode selection.
@@ 7 Conclusion
In this paper, we propose ORBIT, a fully on-policy training framework for learning controllable reasoning models that operates at a target reasoning level while maintaining high reasoning density within each mode. We employ a simple truncation reward and a one-time Expansion–Compression Loop to explore the reasoning frontier, followed by On-Policy Distillation to consolidate multiple modes into a unified model initialized from merged frontier policies, which enables inference-time mode control via natural language prompts. Experiments across LRMs of different sizes show that ORBIT enables controllable reasoning across modes while maintaining competitive accuracy–length trade-offs within each mode compared to existing RL-based efficient reasoning methods.
@@
We elucidate the limitations of this work as follows: First, long-form RL-based training can be computationally expensive for larger models, since each rollout requires autoregressive sampling, which may limit scalability to very large architectures or long-context reasoning tasks. Moreover, when the model output entropy is low, on-policy RL training may suffer from low sample efficiency, limiting further compression progress. Second, our evaluation mainly focuses on reasoning tasks with relatively structured verification signals, and we have not extended experiments to tasks where correctness is difficult to verify with rule-based criteria. Finally, despite the empirical results, we do not establish theoretical guarantees on the convergence or optimality of the proposed training procedure.
@@
Appendix A Intuition for Logarithmic Compression Scheduling
The logarithmic schedule used in the expansion–compression loop can be motivated by a simple structural property under heavy-tailed trajectory length distributions in LLM responses (Jiao et al., 2025).
Let $T:=|o|$ denote the trajectory length. Define the tail probability
$S(L):=\mathbb{P}(T>L),$
which measures the probability mass of trajectories whose length exceeds the context budget $L$.
Over the relevant range of budgets, we approximate the tail distribution by a heavy-tailed form:
$S(L)=CL^{-\alpha},\qquad C>0,\ \alpha>0.$ (13)
Under a multiplicative compression schedule
$L_{k+1}=\rho L_{k},\qquad 0<\rho<1,$ (14)
the truncated trajectory mass satisfies
$\frac{S(L_{k+1})}{S(L_{k})}$ $=\frac{C(\rho L_{k})^{-\alpha}}{CL_{k}^{-\alpha}}=\rho^{-\alpha}.$ (15)
which is independent of $k$. Thus each compression stage imposes a constant relative truncation pressure on the trajectory distribution.
By contrast, for example, under an additive schedule
$L_{k+1}=L_{k}-\Delta,\qquad\Delta>0,$ (16)
the corresponding change becomes
$\frac{S(L_{k+1})}{S(L_{k})}=\left(\frac{L_{k}}{L_{k}-\Delta}\right)^{\alpha},$ (17)
which depends on $L_{k}$ and grows as $L_{k}$ decreases.
We therefore take a logarithmic scheduling with $\rho=0.5$ in ORBIT for a balance between stable truncation and clear separation of different reasoning policies.
Appendix B Mode-aware initialization via Model Merging
As described in Sec. 3.2, optimizing Eq. 12 in practice faces a "cold-start" challenge. To mitigate this misalignment, we proposed using model merging as an initialization of the student model.
Model merging acts as a simple linear combination of model parameters. Numerous methods have been proposed, we use average merging (Wortsman et al., 2022) here for simplicity. This process can be defined as follows:
$\phi=\frac{1}{K}\sum_{k=1}^{K}\theta_{k}$ (18)
We compare two initialization strategies: the merged model and the final-stage policy from sequential multi-stage RL, as these are the only models trained under all reasoning modes.
We first use the Accumulated Misalignment Rate (AMR), which measures whether the top-1 token probabilities differ between two models. This metric aligns naturally with the on-policy distillation process and provides an intuitive visualization of behavioral differences. Specifically, let $a_{t}=\arg\max\pi(o_{t}|q,o_{<t})$ and $\hat{a}_{t}=\arg\max\pi_{\theta_{k}}(o_{t}|q,o_{<t})$. The AMR is defined as:
$\begin{split}\text{AMR}(\pi,\pi_{\theta_{k}})&=\mathbb{E}_{q\sim\mathcal{D},o\sim\pi(\cdot|q)}\left[\frac{1}{|o|}\sum_{t=1}^{|o|}\mathbb{I}(a_{t}\neq\hat{a}_{t})\right]\end{split}$ (19)
Figure 6: Misalignment rate on sequences generated by the merged model and the low-mode model.
Figure 7: RKL training loss dynamics under different initializations in OPD. 30 steps denotes convergence.
As shown in Figure 6, this merged initialization significantly reduces the initial misalignment gap, providing a more stable starting point and accelerating the convergence of subsequent RKL distillation.
Appendix C Training Details
Training Configuration. We maintain consistent hyperparameters throughout the exploration RL stage, using a batch size of $256$, rejection pools of $512$, and a learning rate of $1\times 10^{-6}$ without warm-up. For generation, we apply standard rollout settings: temperature $1$, top-p $1$, top-k 1-1, and $8$ rollouts per problem. We train DeepSeek-Distill-Qwen-1.5B using the DAPO-Math-17K dataset (Yu et al., 2025), while Qwen3-4B-2507-Thinking and Nemotron-7B are trained on the Polaris-53K (An et al., 2025).
For exploitation, we train OPD with a batch size of $1024$, single generation per prompt, and a token-wise KL penalty coefficient of $1.0$.
Appendix D System Prompts For Each Reasoning Mode
@@
To build a semantically meaningful guidance for discrete mode separated reasoning behavior, we follow Yang et al. (2025b) and generalized it to four modes, where each one of them represents semactically aligned expected reasoning behavior under corresponding mode.
@@
/* Low Mode */ You have extremely limited time to think and respond to the user’s query. Every additional second of processing and reasoning incurs a significant resource cost, which could affect efficiency and effectiveness. Your task is to prioritize speed without sacrificing essential clarity or accuracy. Provide the most direct and concise answer possible. Avoid unnecessary steps, reflections, verification, or refinements UNLESS ABSOLUTELY NECESSARY. Your primary goal is to deliver a quick, clear and correct response. /* Mid Mode */ You have sufficient time to think and respond to the user’s query, allowing for a more thoughtful and in-depth answer. However, be aware that the longer you take to reason and process, the greater the associated resource costs and potential consequences. While you should not rush, aim to balance the depth of your reasoning with efficiency. Prioritize providing a well-thought-out response, but do not overextend your thinking if the answer can be provided with a reasonable level of analysis. Use your reasoning time wisely, focusing on what is essential for delivering an accurate response without unnecessary delays and overthinking. /* High Mode */ You have a substantial amount of time to think and respond. The pressure on reasoning cost is low, allowing you to prioritize accuracy and robustness over speed. You should verify your key steps and consider alternative interpretations to ensure your answer is reliable. While you are encouraged to think deeply, aim for a confident solution rather than an endless exploration. If you encounter ambiguity, take the time to resolve it, but remain focused on deriving the correct conclusion efficiently within a generous budget. /* Xhigh Mode */ You have unlimited time to think and respond. There are absolutely no constraints on reasoning length or associated costs. Your singular goal is to achieve the highest possible level of accuracy and certainty. You must exhaustively explore the problem space, considering all edge cases, theoretical nuances, and alternative methodologies. Engage in rigorous, iterative self-correction and deep verification to eliminate even the slightest possibility of error. Do not settle for a ’likely’ correct answer; explore every angle until you are certain. Perfection is the only standard.
@@
Table 3: System prompts used for triggering Low, Mid, High, Xhigh modes
Appendix E Additional Evaluation Results
@@
- Teachers Low 28.3 2.6 19.0 2.0 11.0 2.0 34.3 1.8 37.6 1.4 26.0 2.0
- Teachers Mid 35.2 4.9 26.4 4.5 10.4 3.6 32.9 2.9 38.2 2.2 28.6 3.6
- Teachers High 40.0 7.0 29.8 6.7 14.0 6.3 39.8 4.8 40.7 3.8 32.9 5.7
- Teachers Xhigh 42.5 10.7 32.9 11.0 16.0 11.7 39.5 8.3 40.0 7.7 34.2 9.9
@@
- Teachers Low 64.4 5.0 54.3 5.5 36.6 5.2 57.3 1.9 71.5 0.9 56.8 3.7
- Teachers Mid 66.3 5.9 57.7 6.4 40.0 6.4 63.8 3.1 72.3 1.9 60.0 4.7
- Teachers High 75.1 9.4 69.1 10.6 46.0 11.7 65.4 4.9 73.4 3.0 65.8 7.9
- Teachers Xhigh 83.2 16.8 78.2 19.0 52.7 22.7 67.1 7.9 74.1 5.8 71.1 14.4
@@
- Teachers Low 55.2 4.5 33.1 4.5 29.8 3.7 29.3 2.2 39.2 1.7 37.3 3.3
- Teachers Mid 67.1 6.7 44.5 7.6 36.3 6.6 32.4 4.4 41.2 3.0 44.3 5.7
- Teachers High 74.3 11.1 62.1 12.9 41.9 12.1 30.9 7.9 42.6 5.3 50.4 9.9
- Teachers Xhigh 75.9 13.2 65.1 14.9 43.8 15.0 33.7 9.8 43.1 6.5 52.3 11.9
Table 4: Comparison of reasoning modes of Teachers across various reasoning tasks.
Setting AIME 24 AIME 25 Beyond AIME GPQA MMLU Overall
avg@32 Toks avg@32 Toks avg@8 Toks avg@8 Toks avg@1 Toks Avg. Toks
DeepSeek-Distill-Qwen-1.5B
Prompt Steering Low 26.1 15.1 21.9 14.2 8.5 13.8 28.7 8.4 34.5 5.4 23.9 11.4
Prompt Steering Mid 26.8 15.5 22.6 14.9 10.0 15.0 28.6 9.9 35.3 6.3 24.7 12.3
Prompt Steering High 28.4 16.2 22.5 15.9 10.5 16.2 28.9 11.4 35.3 6.8 25.1 13.3
Prompt Steering Xhigh 28.2 16.1 20.0 15.7 10.0 15.9 29.0 12.3 35.1 7.0 24.5 13.4
Qwen3-4B-2507-Thinking
Prompt Steering Low 79.3 16.6 77.0 17.1 47.1 20.4 63.6 7.3 73.1 3.8 68.0 13.0
Prompt Steering Mid 79.0 19.9 77.3 20.4 54.4 26.8 66.7 8.5 72.8 4.6 70.0 16.0
Prompt Steering High 81.1 20.4 77.7 20.8 53.5 27.7 65.4 9.1 74.1 5.0 70.4 16.6
Prompt Steering Xhigh 81.9 21.1 79.1 21.9 53.6 28.8 66.9 9.9 73.7 6.2 71.0 17.6
Openmath-Nemotron-7B
Prompt Steering Low 70.8 11.3 57.5 13.7 41.3 13.1 31.9 8.2 42.6 5.5 48.8 10.4
Prompt Steering Mid 70.0 11.8 57.5 13.1 40.3 13.6 33.4 8.2 42.4 5.7 48.7 10.5
Prompt Steering High 73.3 11.1 55.8 13.3 41.8 13.5 34.4 8.6 42.5 5.7 49.6 10.4
Prompt Steering Xhigh 69.1 12.1 61.0 13.1 42.6 13.3 34.1 8.5 42.3 5.8 49.8 10.6
Table 5: Prompt steering results on initial models across various reasoning tasks.
E.1 Frontier Teacher Model Performance